# Recursive PAC-Bayes: A Frequentist Approach to Sequential Prior Updates

### v1

@@ 1 Introduction
Selection of a “good” prior plays an important role in the PAC-Bayesian bounds. If one manages to foresee which prediction rules are likely to produce low prediction error and allocate a higher prior mass for them, then the bounds are tighter, because the posterior only needs to make a small deviation from the prior. But if the prior mass on well-performing prediction rules is small, the bounds are loose. A major technique to design good priors is to use part of the data to estimate a good prior and the rest of the data to compute a PAC-Bayes bound. It is known as data-dependent or data-informed priors (Ambroladze et al., 2007). However, all existing approaches to data-informed priors have three major disadvantages. The first is that the bounds are computed on “the rest of the data” that were not used in construction of the prior. Thus, the sample size in the bounds is only a fraction of the total sample size. Therefore, empirically data-informed priors are not always helpful. In many cases starting with an uninformed prior and using all the data to compute the posterior and the bound turns to be superior to sacrificing part of the data for prior construction (Ambroladze et al., 2007, Mhammedi et al., 2019). The second disadvantage is that all the confidence information about the prior is lost in the process. In particular, a prior trained on a few data points is treated in the same way as a prior trained on a lot of data. And a third related disadvantage is that sequential data processing is detrimental, because the bounds only depend on the size of the last chunk and all the confidence information from processing earlier chunks is lost in the process.
@@
We note that while several recent papers experimented with sequential posterior updates by using martingale-style analysis, in all these works the prior remained fixed and only the posterior was changing (Chugg et al., 2023, Biggs and Guedj, 2023, Rodríguez-Gálvez et al., 2024). Another line of work used tools from online learning to derive PAC-Bayesian bounds (Jang et al., 2023), and in this context Haddouche and Guedj (2023) have used sequential prior updates, but their bounds hold for a uniform aggregation of sequentially constructed posteriors, which is different from standard posteriors studied in our work. The confidence bounds in their work come primarily from aggregation rather than confidence in individual posteriors in the sequence. Our work is the first one allowing sequential prior updates without loss of confidence information.
@@ 2 The evolution of data-informed priors and the idea of Recursive PAC-Bayes
In this section we briefly survey the evolution of data-informed priors, and then present our construction of Recursive PAC-Bayes. We consider the standard binary classification setting, with $\mathcal{X}$ being a sample space, $\mathcal{Y}$ a label space, $\mathcal{H}$ a set of prediction rules $h:\mathcal{X}\to\mathcal{Y}$, and $\ell(h(X),Y)=\mathds{1}{\left(h(X)\neq Y\right)}$ the zero-one loss function, where $\mathds{1}{\left(\cdot\right)}$ denotes the indicator function. We let $\mathcal{D}$ denote a distribution on $\mathcal{X}\times\mathcal{Y}$ and $S=\left\{(X_{1},Y_{1}),\dots,(X_{n},Y_{n})\right\}$ an i.i.d. sample from $\mathcal{D}$. Let $L(h)=\mathbb{E}_{(X,Y)\sim\mathcal{D}}[\ell(h(X),Y)]$ be the expected and $\hat{L}(h,S)=\frac{1}{n}\sum_{i=1}^{n}\ell(h(X_{i}),Y_{i})$ the empirical loss.
@@
To construct $\pi_{1},\dots,\pi_{T}$ we split the data $S$ into $T$ non-overlapping subsets, $S=S_{1}\cup\cdots\cup S_{T}$. We restrict $\pi_{t}$ to depend on $U^{\mathrm{train}}_{t}=\bigcup_{s=1}^{t}S_{s}$ only, and we use $U^{\mathrm{val}}_{t}=\bigcup_{s=t}^{T}S_{s}$ to estimate (recursively) $\mathbb{E}_{\pi_{t}}[L(h)]$ (see Figure 1 in Appendix A). Note that $S_{t}$ is used both for construction of $\pi_{t}$ and for estimation of $\mathbb{E}_{\pi_{t}}[L(h)]$ (it is both in $U^{\mathrm{train}}_{t}$ and $U^{\mathrm{val}}_{t}$), resulting in efficient use of the data. It is possible to use any standard PAC-Bayes bound, e.g., Theorem 1, to bound $\mathbb{E}_{\pi_{1}}[L(h)]$, and any PAC-Bayes-Empirical-Bernstein-style bound or the PAC-Bayes-split-kl bound to bound the excess losses $\mathbb{E}_{\pi_{t}}[L(h)-\gamma_{t}\mathbb{E}_{\pi_{t-1}}[L(h)]]$. The excess losses take more than three values, so in the next section we present a generalization of the PAC-Bayes-split-kl inequality to general discrete random variables, which may be of independent interest. The Recursive PAC-Bayes bound is presented in Section 4.
@@ 3.2 PAC-Bayes-Split-kl inequality
Let $f:\mathcal{H}\times\mathcal{Z}\to\left\{b_{0},\dots,b_{K}\right\}$ be a $(K+1)$-valued loss function. (To connect it to the earlier examples, in the binary prediction case we would have $\mathcal{Z}=\mathcal{X}\times\mathcal{Y}$ with elements $Z=(X,Y)$ and $f(h,Z)=\ell(h(X),Y)$, but we will need a more general space $\mathcal{Z}$ later.) For $j\in\left\{1,\dots,K\right\}$ let $f_{j}(\cdot,\cdot)=\mathds{1}{\left(f(\cdot,\cdot)\geq b_{j}\right)}$. Let $\mathcal{D}_{Z}$ be an unknown distribution on $\mathcal{Z}$. For $h\in\mathcal{H}$ let $F(h)=\mathbb{E}_{\mathcal{D}_{Z}}[f(h,Z)]$ and $F_{j}(h)=\mathbb{E}_{\mathcal{D}_{Z}}[f_{j}(h,Z)]$. Let $S=\left\{Z_{1},\dots,Z_{n}\right\}$ be an i.i.d. sample according to $\mathcal{D}_{Z}$ and $\hat{F}_{j}(h,S)=\frac{1}{n}\sum_{i=1}^{n}f_{j}(h,Z_{i})$.
@@
$\mathbb{P}{\left(\exists\rho\in\mathcal{P}:\mathbb{E}_{\rho}[F(h)]\geq b_{0}+\sum_{j=1}^{K}\alpha_{j}\operatorname{kl}^{-1,+}\left(\mathbb{E}_{\rho}[\hat{F}_{j}(h,S)],\frac{\operatorname{KL}(\rho\|\pi)+\ln\frac{2K\sqrt{n}}{\delta}}{n}\right)\right)}\leq\delta,$
@@
We have $f(\cdot,\cdot)=b_{0}+\sum_{j=1}^{K}\alpha_{j}f_{j}(\cdot,\cdot)$ and $F(h)=b_{0}+\sum_{j=1}^{K}\alpha_{j}F_{j}(h)$. Therefore,
$\mathbb{P}{\left(\exists\rho\in\mathcal{P}:\mathbb{E}_{\rho}[F(h)]\geq b_{0}+\sum_{j=1}^{K}\alpha_{j}\operatorname{kl}^{-1,+}\left(\mathbb{E}_{\rho}[\hat{F}_{j}(h,S)],\frac{\operatorname{KL}(\rho\|\pi)+\ln\frac{2K\sqrt{n}}{\delta}}{n}\right)\right)}\\ \leq\mathbb{P}{\left(\exists\rho\in\mathcal{P}\text{ and }\exists j:\mathbb{E}_{\rho}[F_{j}(h)]\geq\operatorname{kl}^{-1,+}\left(\mathbb{E}_{\rho}[\hat{F}_{j}(h,S)],\frac{\operatorname{KL}(\rho\|\pi)+\ln\frac{2K\sqrt{n}}{\delta}}{n}\right)\right)}\leq\delta,$
where the first inequality is by the decomposition of $F$ and the second inequality is by the union bound and application of Theorem 1 to $F_{j}$ (note that $f_{j}$ is a zero-one loss function). ∎
@@ 4 Recursive PAC-Bayes bound
Now we derive a Recursive PAC-Bayes bound based on the loss decomposition in equation (2). In order to bound $\mathbb{E}_{\pi_{t}}[L(h)-\gamma_{t}\mathbb{E}_{\pi_{t-1}}[L(h^{\prime})]]$ we need empirical estimates of $L(h)-\gamma_{t}\mathbb{E}_{\pi_{t-1}}[L(h^{\prime})]$. We denote $F_{\gamma_{t}}(h,\pi_{t-1})=L(h)-\gamma_{t}\mathbb{E}_{\pi_{t-1}}[L(h^{\prime})]=\mathbb{E}_{\mathcal{D}\times\pi_{t-1}}[\ell(h(X),Y)-\gamma_{t}\ell(h^{\prime}(X),Y)]$, where $\mathcal{D}\times\pi_{t-1}$ is a product distribution on $\mathcal{X}\times\mathcal{Y}\times\mathcal{H}$ and $h^{\prime}\in\mathcal{H}$ is sampled according to $\pi_{t-1}$. We define $f_{\gamma_{t}}(h,(X,Y,h^{\prime}))=\ell(h(X),Y)-\gamma_{t}\ell(h^{\prime}(X),Y)\in\left\{-\gamma_{t},0,1-\gamma_{t},1\right\}$, then $F_{\gamma_{t}}(h,\pi_{t-1})=\mathbb{E}\left[f_{\gamma_{t}}(h,(X,Y,h^{\prime}))\right]$. We let $\left\{b_{t|0},b_{t|1},b_{t|2},b_{t|3}\right\}=\left\{-\gamma_{t},0,1-\gamma_{t},1\right\}$ and $f_{\gamma_{t}|j}(h,(X,Y,h^{\prime}))=\mathds{1}{\left(f_{\gamma_{t}}(h,(X,Y,h^{\prime}))\geq b_{t|j}\right)}$. We let $F_{\gamma_{t}|j}(h,\pi_{t-1})=\mathbb{E}\left[f_{\gamma_{t}|j}(h,(X,Y,h^{\prime}))\right]$, then $F_{\gamma_{t}}(h,\pi_{t-1})=-\gamma_{t}+\sum_{j=1}^{3}(b_{t|j}-b_{t|j-1})F_{\gamma_{t}|j}(h,\pi_{t-1})$.
For each sample $(X_{i},Y_{i})$ we sample a prediction rule $h_{i}$ to serve for $h^{\prime}$. We let $\hat{\pi}_{t-1}=\left\{h_{t-1,1},h_{t-1,2},\dots\right\}$ to be the resulting sequence of prediction rules sampled independently according to $\pi_{t-1}$. We only use the first elements of the sequence in accordance with the size of the sample used in the corresponding estimation. We define $\hat{F}_{\gamma_{t}|j}(h,U^{\mathrm{val}}_{t},\hat{\pi}_{t-1})=\frac{1}{n^{\mathrm{val}}_{t}}\sum_{(X,Y)\in U^{\mathrm{val}}_{t},h^{\prime}\in\hat{\pi}_{t-1}}f_{\gamma_{t}|j}(h,(X,Y,h^{\prime}))$, where $n^{\mathrm{val}}_{t}=|U^{\mathrm{val}}_{t}|$ and we take the first $n^{\mathrm{val}}_{t}$ elements from $\hat{\pi}_{t-1}$, one for each sample. Note that $\mathbb{E}\left[\hat{F}_{\gamma_{t}|j}(h,S,\hat{\pi}_{t-1})\right]=F_{\gamma_{t}|j}(h,\pi_{t-1})$. Now we are ready to state the bound.
@@
$\mathcal{E}_{t}(\pi_{t},\gamma_{t})=-\gamma_{t}+\sum_{j=1}^{3}(b_{t|j}-b_{t|j-1})\operatorname{kl}^{-1,+}\left(\mathbb{E}_{\pi_{t}}\left[\hat{F}_{\gamma_{t}|j}(h,U^{\mathrm{val}}_{t},\hat{\pi}_{t-1}^{*})\right],\frac{\operatorname{KL}(\pi_{t}\|\pi_{t-1}^{*})+\ln\frac{6T\sqrt{n^{\mathrm{val}}_{t}}}{\delta}}{n^{\mathrm{val}}_{t}}\right)$
@@
$\operatorname{B}_{t}(\pi_{t})=\mathcal{E}_{t}(\pi_{t})+\gamma_{t}\operatorname{B}_{t-1}(\pi_{t-1}^{*}).$ (3)

### v3

@@ 1 Introduction
Selection of a “good” prior plays an important role in the PAC-Bayesian bounds. If one manages to foresee which prediction rules are likely to produce low prediction error and allocate a higher prior mass for them, then the bounds are tighter, because the posterior only needs to make a small deviation from the prior. But if the prior mass on well-performing prediction rules is small, the bounds are loose. A major technique to design good priors is to use part of the data to estimate a good prior and the rest of the data to compute a PAC-Bayes bound. It is known as data-dependent or data-informed priors (Ambroladze et al., 2007). However, all existing approaches to data-informed priors have three major disadvantages. The first is that the bounds are computed on “the rest of the data” that were not used in construction of the prior. Thus, the sample size in the bounds is only a fraction of the total sample size. Therefore, empirically data-informed priors are not always helpful. In many cases starting with an uninformed prior and using all the data to compute the posterior and the bound turns to be superior to sacrificing part of the data for prior construction (Ambroladze et al., 2007, Mhammedi et al., 2019). The second disadvantage is that all the confidence information about the prior is lost in the process. In particular, a prior trained on a few data points is treated in the same way as a prior trained on a lot of data. And a third related disadvantage is that sequential data processing provides no benefit, because the bounds only depend on the size of the last chunk and all the confidence information from processing earlier chunks is lost in the process.
@@
We note that while several recent papers experimented with sequential posterior updates by using martingale-style analysis, in all these works the prior remained fixed and only the posterior was changing (Chugg et al., 2023, Biggs and Guedj, 2023, Rodríguez-Gálvez et al., 2024). The work on sequential posterior updates is orthogonal to our contribution and can be combined with it. Namely, it is possible to apply sequential posterior updates in-between sequential prior updates. Another line of work used tools from online learning to derive PAC-Bayesian bounds (Jang et al., 2023), and in this context Haddouche and Guedj (2023) have used sequential prior updates, but their bounds hold for a uniform aggregation of sequentially constructed posteriors, which is different from standard posteriors studied in our work. The confidence bounds in their work come primarily from aggregation rather than confidence in individual posteriors in the sequence (the denominator of their bounds depends on the number of aggregated posteriors). The need to construct and maintain a large number of posteriors has a negative impact on the computational efficiency. Our work is the first one allowing sequential prior updates without loss of confidence information.
@@ 2 The evolution of data-informed priors and the idea of Recursive PAC-Bayes
In this section we briefly survey the evolution of data-informed priors, and then present our construction of Recursive PAC-Bayes. We consider the standard classification setting, with $\mathcal{X}$ being a sample space, $\mathcal{Y}$ a label space, $\mathcal{H}$ a set of prediction rules $h:\mathcal{X}\to\mathcal{Y}$, and $\ell(h(X),Y)=\mathds{1}{\left(h(X)\neq Y\right)}$ the zero-one loss function, where $\mathds{1}{\left(\cdot\right)}$ denotes the indicator function. We let $\mathcal{D}$ denote a distribution on $\mathcal{X}\times\mathcal{Y}$ and $S=\left\{(X_{1},Y_{1}),\dots,(X_{n},Y_{n})\right\}$ an i.i.d. sample from $\mathcal{D}$. Let $L(h)=\mathbb{E}_{(X,Y)\sim\mathcal{D}}[\ell(h(X),Y)]$ be the expected and $\hat{L}(h,S)=\frac{1}{n}\sum_{i=1}^{n}\ell(h(X_{i}),Y_{i})$ the empirical loss.
@@
To construct $\pi_{1},\dots,\pi_{T}$ we split the data $S$ into $T$ non-overlapping subsets, $S=S_{1}\cup\cdots\cup S_{T}$. We restrict $\pi_{t}$ to depend on $U^{\mathrm{train}}_{t}=\bigcup_{s=1}^{t}S_{s}$ only, and we use $U^{\mathrm{val}}_{t}=\bigcup_{s=t}^{T}S_{s}$ to estimate (recursively) $\mathbb{E}_{\pi_{t}}[L(h)]$ (see Figures 1 and 2 in Appendix A for a graphical illustration). Note that $S_{t}$ is used both for construction of $\pi_{t}$ and for estimation of $\mathbb{E}_{\pi_{t}}[L(h)]$ (it is both in $U^{\mathrm{train}}_{t}$ and $U^{\mathrm{val}}_{t}$), resulting in efficient use of the data. It is possible to use any standard PAC-Bayes bound, e.g., Theorem 1, to bound $\mathbb{E}_{\pi_{1}}[L(h)]$, and any PAC-Bayes-Empirical-Bernstein-style bound or the PAC-Bayes-split-kl bound to bound the excess losses $\mathbb{E}_{\pi_{t}}[L(h)-\gamma_{t}\mathbb{E}_{\pi_{t-1}}[L(h)]]$. The excess losses take more than three values, so in the next section we present a generalization of the PAC-Bayes-split-kl inequality to general discrete random variables, which may be of independent interest. The Recursive PAC-Bayes bound is presented in Section 4.
@@ 3.2 PAC-Bayes-Split-kl inequality
Let $f:\mathcal{H}\times\mathcal{Z}\to\left\{b_{0},\dots,b_{K}\right\}$ be a $(K+1)$-valued loss function. (To connect it to the earlier examples, in the binary prediction case we would have $\mathcal{Z}=\mathcal{X}\times\mathcal{Y}$ with elements $Z=(X,Y)$ and $f(h,Z)=\ell(h(X),Y)$, but we will need a more general space $\mathcal{Z}$ later.) For $j\in\left\{1,\dots,K\right\}$ let $f_{|j}(\cdot,\cdot)=\mathds{1}{\left(f(\cdot,\cdot)\geq b_{j}\right)}$. Let $\mathcal{D}_{Z}$ be an unknown distribution on $\mathcal{Z}$. For $h\in\mathcal{H}$ let $F(h)=\mathbb{E}_{\mathcal{D}_{Z}}[f(h,Z)]$ and $F_{|j}(h)=\mathbb{E}_{\mathcal{D}_{Z}}[f_{|j}(h,Z)]$. Let $S=\left\{Z_{1},\dots,Z_{n}\right\}$ be an i.i.d. sample according to $\mathcal{D}_{Z}$ and $\hat{F}_{|j}(h,S)=\frac{1}{n}\sum_{i=1}^{n}f_{|j}(h,Z_{i})$.
@@
$\mathbb{P}{\left(\exists\rho\in\mathcal{P}:\mathbb{E}_{\rho}[F(h)]\geq b_{0}+\sum_{j=1}^{K}\alpha_{j}\operatorname{kl}^{-1,+}\left(\mathbb{E}_{\rho}[\hat{F}_{|j}(h,S)],\frac{\operatorname{KL}(\rho\|\pi)+\ln\frac{2K\sqrt{n}}{\delta}}{n}\right)\right)}\leq\delta,$
@@
We have $f(\cdot,\cdot)=b_{0}+\sum_{j=1}^{K}\alpha_{j}f_{|j}(\cdot,\cdot)$ and $F(h)=b_{0}+\sum_{j=1}^{K}\alpha_{j}F_{|j}(h)$. Therefore,
$\mathbb{P}{\left(\exists\rho\in\mathcal{P}:\mathbb{E}_{\rho}[F(h)]\geq b_{0}+\sum_{j=1}^{K}\alpha_{j}\operatorname{kl}^{-1,+}\left(\mathbb{E}_{\rho}[\hat{F}_{|j}(h,S)],\frac{\operatorname{KL}(\rho\|\pi)+\ln\frac{2K\sqrt{n}}{\delta}}{n}\right)\right)}\\ \leq\mathbb{P}{\left(\exists\rho\in\mathcal{P}\text{ and }\exists j:\mathbb{E}_{\rho}[F_{|j}(h)]\geq\operatorname{kl}^{-1,+}\left(\mathbb{E}_{\rho}[\hat{F}_{|j}(h,S)],\frac{\operatorname{KL}(\rho\|\pi)+\ln\frac{2K\sqrt{n}}{\delta}}{n}\right)\right)}\leq\delta,$
where the first inequality is by the decomposition of $F$ and the second inequality is by the union bound and application of Theorem 1 to $F_{|j}$ (note that $f_{|j}$ is a zero-one loss function). ∎
@@ 4 Recursive PAC-Bayes bound
Now we derive a Recursive PAC-Bayes bound based on the loss decomposition in equation (2). We aim to bound $\mathbb{E}_{\pi_{t}}[L(h)-\gamma_{t}\mathbb{E}_{\pi_{t-1}}[L(h^{\prime})]]$, which we denote by
$F_{\gamma_{t},\pi_{t-1}}(h)=L(h)-\gamma_{t}\mathbb{E}_{\pi_{t-1}}[L(h^{\prime})]=\mathbb{E}_{\mathcal{D}\times\pi_{t-1}}[\ell(h(X),Y)-\gamma_{t}\ell(h^{\prime}(X),Y)],$
where $\mathcal{D}\times\pi_{t-1}$ is a product distribution on $\mathcal{X}\times\mathcal{Y}\times\mathcal{H}$ and $h^{\prime}\in\mathcal{H}$ is sampled according to $\pi_{t-1}$. We further define
$f_{\gamma_{t}}(h,(X,Y,h^{\prime}))=\ell(h(X),Y)-\gamma_{t}\ell(h^{\prime}(X),Y)\in\left\{-\gamma_{t},0,1-\gamma_{t},1\right\},$
then $F_{\gamma_{t},\pi_{t-1}}(h)=\mathbb{E}_{\mathcal{D}\times\pi_{t-1}}[f_{\gamma_{t}}(h,(X,Y,h^{\prime}))]$. In order to apply Theorem 4, we represent $f_{\gamma_{t}}$ as a superposition of binary functions. For this purpose we let $\left\{b_{t|0},b_{t|1},b_{t|2},b_{t|3}\right\}=\left\{-\gamma_{t},0,1-\gamma_{t},1\right\}$ and define $f_{\gamma_{t}|j}(h,(X,Y,h^{\prime}))=\mathds{1}{\left(f_{\gamma_{t}}(h,(X,Y,h^{\prime}))\geq b_{t|j}\right)}$. We let $F_{\gamma_{t},\pi_{t-1}|j}(h)=\mathbb{E}_{\mathcal{D}\times\pi_{t-1}}[f_{\gamma_{t}|j}(h,(X,Y,h^{\prime}))]$, then $F_{\gamma_{t},\pi_{t-1}}(h)=-\gamma_{t}+\sum_{j=1}^{3}(b_{t|j}-b_{t|j-1})F_{\gamma_{t},\pi_{t-1}|j}(h)$.
Now we construct an empirical estimate of $F_{\gamma_{t},\pi_{t-1}|j}(h)$. We first let $\hat{\pi}_{t-1}=\left\{h^{\pi_{t-1}}_{1},h^{\pi_{t-1}}_{2},\dots\right\}$ be a sequence of prediction rules sampled independently according to $\pi_{t-1}$. We define ${U^{\mathrm{val}}_{t}\circ\hat{\pi}_{t-1}}=\left\{\left(X_{i},Y_{i},h^{\pi_{t-1}}_{i}\right):(X_{i},Y_{i})\in U^{\mathrm{val}}_{t}\right\}$. In words, for every sample $(X_{i},Y_{i})\in U^{\mathrm{val}}_{t}$ we sample a prediction rule $h^{\pi_{t-1}}_{i}$ according to $\pi_{t-1}$ and put the triplet $(X_{i},Y_{i},h^{\pi_{t-1}}_{i})$ in $U^{\mathrm{val}}_{t}\circ\hat{\pi}_{t-1}$. The triplets $(X_{i},Y_{i},h^{\pi_{t-1}}_{i})$ correspond to the random variables $Z$ in Theorem 4. We note that $|U^{\mathrm{val}}_{t}|=|U^{\mathrm{val}}_{t}\circ\hat{\pi}_{t-1}|$, and we let $n^{\mathrm{val}}_{t}=|U^{\mathrm{val}}_{t}|$. We define the empirical estimate of $F_{\gamma_{t},\pi_{t-1}|j}(h)$ as ${\hat{F}_{\gamma_{t}|j}(h,U^{\mathrm{val}}_{t}\circ\hat{\pi}_{t-1})}=\frac{1}{n^{\mathrm{val}}_{t}}\sum_{(X,Y,h^{\prime})\in U^{\mathrm{val}}_{t}\circ\hat{\pi}_{t-1}}f_{\gamma_{t}|j}(h,(X,Y,h^{\prime}))$. Note that $\mathbb{E}_{\mathcal{D}\times\pi_{t-1}}[\hat{F}_{\gamma_{t}|j}(h,U^{\mathrm{val}}_{t}\circ\hat{\pi}_{t-1})]=F_{\gamma_{t},\pi_{t-1}|j}(h)$, therefore, we can use Theorem 4 to bound $\mathbb{E}_{\pi_{t}}[F_{\gamma_{t},\pi_{t-1}}(h)]$ using its empirical estimates. We are now ready to state the bound.
@@
$\mathcal{E}_{t}(\pi_{t},\gamma_{t})=-\gamma_{t}+\sum_{j=1}^{3}(b_{t|j}-b_{t|j-1})\operatorname{kl}^{-1,+}\left(\mathbb{E}_{\pi_{t}}\left[\hat{F}_{\gamma_{t}|j}(h,U^{\mathrm{val}}_{t}\circ\hat{\pi}_{t-1}^{*})\right],\frac{\operatorname{KL}(\pi_{t}\|\pi_{t-1}^{*})+\ln\frac{6T\sqrt{n^{\mathrm{val}}_{t}}}{\delta}}{n^{\mathrm{val}}_{t}}\right)$
@@
$\operatorname{B}_{t}(\pi_{t})=\mathcal{E}_{t}(\pi_{t},\gamma_{t})+\gamma_{t}\operatorname{B}_{t-1}(\pi_{t-1}^{*}).$ (3)